OpenStudy (anonymous):

Help with functions? f(x)=square root of x-2 g(x)=x^2 find (f*g)(x). would it just be square root of x-2 * x^2? How would I go about multiplying?

ganeshie8 (ganeshie8):

thats right !

ganeshie8 (ganeshie8):

(f*g)(x) = f(x) * g(x) = (x-2)*x^2

ganeshie8 (ganeshie8):

first multiply x wid x^2, then multiply -2 wid x^2

OpenStudy (anonymous):

how would you multiply the square root of (x-2) * x^2 in the function?

OpenStudy (anonymous):

oh, i see

ganeshie8 (ganeshie8):

oh u have square root, wait a sec

OpenStudy (anonymous):

yes, so it;s the square root of x-2

ganeshie8 (ganeshie8):

(f*g)(x) = f(x) * g(x) = \(\sqrt{x-2}*x^2\)

ganeshie8 (ganeshie8):

= \(x^2\sqrt{x-2}\)

ganeshie8 (ganeshie8):

leave it like that, we cannot simplify it any further.

OpenStudy (anonymous):

great, thanks for your help!